### v1

@@ 2.3 Observability and reconstructability
(20) $z_{0}=A^{n-1}\Lambda_{\mathrm{z},n}(y^{0}_{-n+1}):=\Pi_{\mathrm{z},n}(y^{0}_{-n+1})$
where $\Pi_{\mathrm{z},n}:\mathbb{R}^{nn_{\mathrm{y}}}\rightarrow\mathbb{R}^{n_{\mathrm{z}}}$ is the Koopman reconstructability map for autonomous systems. Note that the compliance constraint $\Psi$ is part of $\Lambda_{\mathrm{z},n}$. This gives a different point of view on reconstructability than in the work [46], where the observability notions are discussed based on an explicit definition of the lifting map, i.e., a given selection of the observables $\Phi$. Note that, for a nonlinear system representation with $n_{\mathrm{x}}$ states and an explicit dictionary $\Phi$, the construction $z_{0}=\Phi(x_{0})=\Phi(\Pi_{\mathrm{x},n}(y^{0}_{-n+1}))$ allows to compute the initial lifted $z_{0}$ based on $x_{0}$, using a much smaller amount of lags (at max $n=n_{\mathrm{x}}$) as, typically, $n_{\mathrm{x}}\ll n_{\mathrm{z}}$. As such, the number of necessary lags $n$ does not depend on the dimensionality of the lifted space, but of the original nonlinear system, which drastically reduces the computational complexity.
It is important to emphasize that the conditions discussed in this subsection guarantee local observability and necessary conditions for the local invertibility of (13) and (17). However, for stronger, global guarantees, [1] describes, albeit for a continuous time systems, that if $f$ is Lipschitz continuous and $h$ has a finite amount of nondegenerate critical points, then $n=2n_{\mathrm{x}}+1$ is sufficient to ensure global existence of the reconstructabiltiy map, which is in line with the results in [49].
@@ 3.1 Data generating system
with $u_{k}\in\mathbb{U}\subseteq\mathbb{R}^{n_{\mathrm{u}}}$ the control input and $x_{k}\in\mathbb{X}\subseteq\mathbb{R}^{n_{\mathrm{x}}}$ the state. The signal $e_{k}$ is the sample-path realization of an i.i.d. white noise process of finite variance, taking values in $\mathbb{E}\rightarrow\mathbb{R}^{n_{\mathrm{x}}}$ and being independent from $u$ in the statistical sense. The functions $f_{\mathrm{d}}:\mathbb{X}\times\mathbb{U}\times\mathbb{E}\rightarrow\mathbb{R}^{n_{\mathrm{x}}}$ and $h:\mathbb{X}\rightarrow\mathbb{R}^{n_{\mathrm{y}}}$ are the state-transition and output functions, respectively. It is assumed that the sets $\mathbb{U}$ and $\mathbb{E}$ are such that $\mathbb{X}$ is forward invariant under $f_{\mathrm{d}}$ and $0\in\mathbb{X}$. The objective is to estimate a Koopman model of the deterministic (process) part of (21). This is done using an input-output data sequence $\mathcal{D}_{N}=\{(u_{k},y_{k})\}^{N}_{k=0}$ collected from the system (21). We define next the model structure that we propose to identify a lifted Koopman form of the system under the control input $u$ and noise process $e$.
@@ 3.2 Koopman model structure
where $g(x_{k},0)=0$ and $d(x_{k},u_{k},0)=0$. This decomposition, which always exists, extends the one discussed in [12] and [46] for the noiseless case. To derive an exact finite dimensional Koopman representation, we make the following assumptions.
@@
There exists a finite dimensional dictionary of observables $\Phi:\mathbb{X}\rightarrow\mathbb{R}^{n_{\mathrm{f}}}$ with $\Phi=[\ \phi_{1}\ \cdots\ \phi_{n_{\mathrm{f}}}\ ]^{\top}$ in an appropriate Banach space $\mathcal{F}$ such that:
(23) $\Phi\circ f(\cdot)\in\mathrm{span}\{\Phi\}\qquad\implies\qquad\Phi\circ f(\cdot)=A\Phi(\cdot),\;\text{with }A\in\mathbb{R}^{n_{\mathrm{f}}\times n_{\mathrm{f}}}.$
@@
The output map $h$ can be exactly represented by the observables $\Phi$ in Assumption 1, in other words:
(24) $h\in\mathrm{span}\{\Phi\}\implies h(x_{k})=C\Phi(x_{k}),\;\text{with }C\in\mathbb{R}^{n_{\mathrm{y}}\times n_{\mathrm{f}}}.$
While Assumption 2 can be easily satisfied (for example by including $h$ in the dictionary of observables), Assumption 1 is more challenging due to the finite dimensionality. While there exist methods for exact finite embedding of polynomial systems [10], [11], methods for polyflows [16], or results in immersion theory [51], the conditions for the existence of an exact embedding of more general classes of nonlinear systems are lacking. Hence, it is currently an open question what are the limitations of Assumption 1. Otherwise, (23) only holds in an approximative sense. In this work, we use Assumptions 1 and 2 to derive the exact finite dimensional Koopman form of nonlinear systems with control input and influenced by process and measurement noise and we use the resulting form as our model structure to be identified. However, our results also give a picture about what is required to have accurate enough models in the approximative case, such that the user can make informed decisions on the complexity vs accuracy tradeoff during the identification process.
@@
$+B_{0}$,$\;B(z_{k}),\;B(z_{k},u_{k})\}$, corresponding to a linear, bilinear, input affine or what we call a general model structure. Similar to the $\mathcal{B}$ matrix, the innovation noise term can be considered with various dependencies: $\mathcal{K}\in\{K,\;\sum_{i=1}^{n_{\mathrm{z}}}K_{i}z_{k,i}+K_{0},\;K(z_{k}),$ $K(z_{k},u_{k}),\;K(z_{k},u_{k},e_{k})\}$. Choosing a suitable structural form of $\mathcal{B}$ and $\mathcal{K}$ corresponds to a model structure selection problem, as in classical system identification.
To use the more complex Koopman models for control (not fully LTI), it is possible to cast the model into a linear parameter-varying (LPV) form (see [12] for the noiseless case) by introducing a so called scheduling variable $p_{k}$ that is required to be measurable/observable from the system. For nonlinear systems described by LPV models, there exists a multitude of convex and computationally efficient control synthesis methods, where the user can also shape performance and achieve global guarantees of stability [30]. To cast (26) into an LPV form, we must first note that, generally, the noise $e_{k}$ is not directly measurable, but thanks to the innovation noise setting of (21), $e_{k}=y_{k}-Cz_{k}$ holds. Then, we can conceptually write the LPV form of the Koopman model (26) as:
@@
where $p_{k}=\mu(z_{k},u_{k},y_{k})$ is a scheduling map and $\mathcal{B}_{\mathrm{z}}$ with $\mathcal{K}_{\mathrm{z}}$ belong to a predefined function class such as affine, polynomial or rational, such that $\mathcal{B}_{\mathrm{z}}\circ\mu=\mathcal{B}$, $\mathcal{K}_{\mathrm{z}}\circ\mu=\mathcal{K}$ [12], [30]. Note that the dependencies of $\mu$ are based on the choice of $\mathcal{B}$ and $\mathcal{K}$.
@@ 3.3 Identification problem
The objective is to introduce a parametrized version of (25) to learn the underlying dependencies together with a lifting map using ANNs. This means identifying the lifted state $z_{k}$, the linear maps $A$, $C$, and the nonlinear maps $B$ and $K$. To this end, we introduce an identification cost function that we chose to be the squared prediction error due to its extended use and success in system identification and its close connection to maximum-likelihood estimators under specific settings [22], [23]. For this, a predictor is needed and we derive it next. As a first step, we exploit $e_{k}=y_{k}-Cz_{k}$ in the assumed innovation form (21) and substitute it in (25a) to obtain:
@@
with $\tilde{K}(z_{k},u_{k},y_{k}):=K(z_{k},u_{k},y_{k}-Cz_{k})$. Then, iterating (25b) forward in time, for $n\geq 1$, we arrive at
@@ 3.4 Subspace encoder
To overcome problem (i), in this section, the estimation of the lifted state $z_{k}$ is considered in terms of an encoder. By exploiting input-output data, we use the reconsutructability concept, discussed in the autonomous case, which we now generalize for (25). Starting with observability, the following map holds based on (30):
@@
where, as in the autonomous case, we have the set of nonlinear constraints $\Psi$. For $n\geq 1$, if $\exists(z_{*},w_{*})\in\mathbb{R}^{n_{\mathrm{z}}}\times\mathbb{R}^{nn_{\mathrm{u}}\times(n+1)n_{\mathrm{y}}\times(n+1)n_{\mathrm{y}}}$ for which the Jacobian $\nabla_{(z_{\ast},w_{\ast})}\mathcal{O}_{\mathrm{z},n}$ has full row rank $n_{\mathrm{z}}$, then there exist open sets $\mathbb{Z}_{0}\subseteq\mathbb{R}^{n_{\mathrm{z}}}$, $\mathbb{U}_{0}\subseteq\mathbb{R}^{n_{\mathrm{u}}}$, $\mathbb{Y}_{0}\subseteq\mathbb{R}^{n_{\mathrm{y}}}$, $\mathbb{E}_{0}\subseteq\mathbb{R}^{n_{\mathrm{y}}}$, corresponding to the neighborhood of $(z_{*},w_{*})$ for which $\mathcal{O}_{\mathrm{z},n}$ is partially invertible and (25) is locally observable on $(\mathbb{Z}_{0},\mathbb{U}_{0},\mathbb{Y}_{0},\mathbb{E}_{0})$ [2, 14, 31]. Note that if the representation is locally observable, then the above condition is satisfied for any $n\geq n_{\mathrm{z}}-1$. By inverting $\mathcal{O}_{\mathrm{z},n}$, we get
@@
where $\Lambda_{\mathrm{z},n}:\mathbb{U}^{n}_{0}\times\mathbb{Y}^{n+1}_{0}\times\mathbb{E}^{n+1}_{0}\rightarrow{\mathbb{R}^{n_{\mathrm{z}}}}$ is the observability map. To determine $z_{k}$ based on past input-output data, we derive
@@
$:=\Pi_{\mathrm{z},n}(u^{k-1}_{k-n},y^{k}_{k-n},e^{k}_{k-n})$
where $\Pi_{\mathrm{z},n}:\mathbb{U}^{n}_{0}\times\mathbb{Y}^{n+1}_{0}\times\mathbb{E}^{n+1}_{0}\rightarrow{\mathbb{R}^{n_{\mathrm{z}}}}$ is the recosntructability map. Note that the noise sequence $e^{k}_{k-n}$ is not directly available in practice, hence to compute $z_{k}$ based on (36), again we can exploit the i.i.d. white noise property of $e_{k}$ to arrive at:
@@
which mapping gives an efficient estimator of $z_{k}$ in the conditional mean sense based on past data with a given lag $n$. In fact, (37) functions as an encoder, mapping from the past data to the lifted state $z_{k}$, i.e., a subspace of the lifted state space. However, an exact calculation of the encoder $\bar{\Pi}_{\mathrm{z},n}$ for a given ANN parametrization of $f_{\theta}$ and $h_{\theta}$ is infeasible in practice, due to the required analytic inversion of $\mathcal{O}_{\mathrm{z},n}$ to get $\Lambda_{\mathrm{z},n}$ and the computation of the conditional expectation of $\bar{\Pi}_{\mathrm{z},n}$ under the unknown probability density function of $e_{k}$. Hence, our objective is to learn $\bar{\Pi}_{\mathrm{z},n}$ directly from the data by introducing a parametrized function $\bar{\Pi}^{\eta}_{\mathrm{z},n}$ with parameters $\eta\in\Upsilon\subseteq\mathbb{R}^{n_{\eta}}$, e.g., using an ANN in the general case, which is co-estimated with $A_{\theta}$, $B_{\theta}$, $C_{\theta}$ and $K_{\theta}$.
Note that, similar to the autonomous case discussed in Section 2.3, we can conceptually show that exploiting the observability and reconstructability properties of the nonlinear system (21), the potentially needed number of lags $n\geq n_{\mathrm{z}}-1$ is greatly reduced. Given a lifting map $\Phi:\mathbb{R}^{n_{\mathrm{x}}}\rightarrow\mathbb{R}^{n_{\mathrm{z}}}$, the state $z_{k}=\Phi(x_{k})$ can be calculated based on the reconstructed $x_{k}$ using a number of lags $n\geq n_{\mathrm{x}}-1$, as detailed in [2], for the reconstructabillity map associated to (21). As such, when computing $\bar{\Pi}^{\eta}_{\mathrm{z},n}$, the number of lags needed to estimate $z_{k}$ is related to the dimension of the underlying nonlinear system, rather than that of the lifted system. It is important to note that, while $n_{\mathrm{x}}-1$ guarantees local invertibility, for global observability guarantees one would need to increase the number of lags $n$. For example, works such as [44] and [45] describe a sufficient condition for reconstruction of $x_{k}$ to be $n\geq 2n_{\mathrm{x}}$, for nonlinear systems with deterministic and stochastic forcing, respectively. Note that, as the last step in (29) is $k+n$ instead of $k+n-1$, we subtract 1 from the given value in [44] and [45], which is $12n_{\mathrm{x}}+$ 1. While (34) is more complex than the maps discussed in [44] and [45], these results can still serve as a guideline when performing experiments and choosing $n$. Furthermore, if $n_{\mathrm{x}}$ is known, although $\mathcal{O}_{\mathrm{z},n}$ may be locally invertible for smaller values of $n$, it is still a safe choice to start with $n\geq n_{\mathrm{x}}-1$, which provides local guarantees for reconstructability.

### v3

@@ 2.3 Observability and reconstructability
(20) $z_{0}=A^{n-1}\Lambda_{\mathrm{z},n}(y^{0}_{-n+1})\mathrel{\mathop{\ordinarycolon}}=\Pi_{\mathrm{z},n}(y^{0}_{-n+1})$
where $\Pi_{\mathrm{z},n}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{nn_{\mathrm{y}}}\rightarrow\mathbb{R}^{n_{\mathrm{z}}}$ is the Koopman reconstructability map for autonomous systems. Note that the compliance constraint $\Psi$ is part of $\Lambda_{\mathrm{z},n}$. This gives a different point of view on reconstructability than in the work [surana_obs], where the observability notions are discussed based on an explicit definition of the lifting map, i.e., a given selection of the observables $\Phi$. Note that, for a nonlinear system representation with $n_{\mathrm{x}}$ states and an explicit dictionary $\Phi$, the construction $z_{0}=\Phi(x_{0})=\Phi(\Pi_{\mathrm{x},n}(y^{0}_{-n+1}))$ allows to compute the initial lifted $z_{0}$ based on $x_{0}$, using a much smaller amount of lags (at max $n=n_{\mathrm{x}}$) as, typically, $n_{\mathrm{x}}\ll n_{\mathrm{z}}$. As such, the number of necessary lags $n$ does not depend on the dimensionality of the lifted space, but of the original nonlinear system, which drastically reduces the computational complexity.
It is important to emphasize that the conditions discussed in this subsection guarantee local observability and necessary conditions for the local invertibility of (13) and (17). However, for stronger, global guarantees, [Jose_subnet] describes, albeit for a continuous time systems, that if $f$ is Lipschitz continuous and $h$ has a finite amount of nondegenerate critical points, then $n=2n_{\mathrm{x}}+1$ is sufficient to ensure global existence of the reconstructabiltiy map, which is in line with the results in [Takens].
@@ 3.1 Data generating system
with $u_{k}\in\mathbb{U}\subseteq\mathbb{R}^{n_{\mathrm{u}}}$ the control input and $x_{k}\in\mathbb{X}\subseteq\mathbb{R}^{n_{\mathrm{x}}}$ the state. The signal $e_{k}$ is the sample-path realization of an i.i.d. white noise process of finite variance, taking values in $\mathbb{E}\subseteq\mathbb{R}^{n_{\mathrm{x}}}$ and being independent from $u$ in the statistical sense. The functions $f_{\mathrm{d}}\mathrel{\mathop{\ordinarycolon}}\mathbb{X}\times\mathbb{U}\times\mathbb{E}\rightarrow\mathbb{R}^{n_{\mathrm{x}}}$ and $h\mathrel{\mathop{\ordinarycolon}}\mathbb{X}\rightarrow\mathbb{R}^{n_{\mathrm{y}}}$ are the state-transition and output functions, respectively. It is assumed that the sets $\mathbb{U}$ and $\mathbb{E}$ are such that $\mathbb{X}$ is forward invariant under $f_{\mathrm{d}}$ and $0\in\mathbb{X}$. The objective is to estimate a Koopman model of the deterministic (process) part of (21). This is done using an input-output data sequence $\mathcal{D}_{N}=\{(u_{k},y_{k})\}^{N}_{k=0}$ collected from the system (21). We define next the model structure that we propose to identify a lifted Koopman form of the system under the control input $u$ and noise process $e$.
@@ 3.2 Koopman model structure
where $g(x_{k},0)=0$ and $d(x_{k},u_{k},0)=0$. This decomposition, which always exists, extends the one discussed in [Aut_Iacob_inputs] and [surana_obs] for the noiseless case. To derive an exact finite dimensional Koopman representation, we make the following assumptions.
@@
There exists a finite dimensional dictionary of observables $\Phi\mathrel{\mathop{\ordinarycolon}}\mathbb{X}\rightarrow\mathbb{R}^{n_{\mathrm{f}}}$ with $\Phi=[\ \phi_{1}\ \cdots\ \phi_{n_{\mathrm{f}}}\ ]^{\top}$ in an appropriate Banach space $\mathcal{F}$ such that $\Phi\circ f(\cdot)\in\mathrm{span}\{\Phi\}$, which implies that
(23) $\Phi\circ f(\cdot)=A\Phi(\cdot)\;\text{ with }A\in\mathbb{R}^{n_{\mathrm{f}}\times n_{\mathrm{f}}}.$
@@
The output map $h$ can be exactly represented by the observables $\Phi$ in Assumption 1, in other words $h\in\mathrm{span}\{\Phi\}$, which implies that
[…]
While Assumption 2 can be easily satisfied (for example by including $h$ in the dictionary of observables), Assumption 1 is more challenging due to the finite dimensionality. While there exist methods for exact finite embedding of polynomial systems [Iacob_poly_system], [Iacob_CT_block_embedding], methods for polyflows [Polyflows], or results in immersion theory [Wang_Immersion], the conditions for the existence of an exact embedding of more general classes of nonlinear systems are lacking. Hence, it is currently an open question what are the limitations of Assumption 1, especially in terms of the findings in [liu-ozay-sontag]. Otherwise, (23) only holds in an approximative sense  [^3] For systems where exact finite dimensional embeddings are not possible, future research may provide a way to combine the consistency result described in Section 4 with a method that quantifies the approximation error when Assumption 1 does not hold. For example, it may be possible to characterize or mitigate the $n$-step reconstructabiltiy error through a reprojection approach as discussed in [Pvangoor].. In this work, we use Assumptions 1 and 2 to derive the exact finite dimensional Koopman form of nonlinear systems with control input and influenced by process and measurement noise and we use the resulting form as our model structure to be identified.
@@
$B_{0}$$B(z_{k}),\;B(z_{k},u_{k})\}$, corresponding to a linear, bilinear, input affine or what we call a general model structure. Similar to the $\mathcal{B}$ matrix, the innovation noise term can be considered with various dependencies: $\mathcal{K}\in\{K,\;\sum_{i=1}^{n_{\mathrm{z}}}K_{i}z_{k,i}+K_{0},\;K(z_{k}),$ $K(z_{k},u_{k}),\;K(z_{k},u_{k},e_{k})\}$. Choosing a suitable structural form of $\mathcal{B}$ and $\mathcal{K}$ corresponds to a model structure selection problem, as in classical system identification.
To use the more complex Koopman models for control (not fully LTI), it is possible to cast the model into a linear parameter-varying (LPV) form (see [Aut_Iacob_inputs] for the noiseless case) by introducing a so called scheduling variable $p_{k}$ that is required to be measurable/observable from the system. For nonlinear systems described by LPV models, there exists a multitude of convex and computationally efficient control synthesis methods, where the user can also shape performance and achieve global guarantees of stability [Mohammadpour]. To cast (26) into an LPV form, we must first note that, generally, the noise $e_{k}$ is not directly measurable, but thanks to the innovation noise setting of (21), $e_{k}=y_{k}-Cz_{k}$ holds. Then, we can conceptually write the LPV form of the Koopman model (26) as:
@@
where $p_{k}=\mu(z_{k},u_{k},y_{k})$ is a scheduling map and $\mathcal{B}_{\mathrm{z}}$ with $\mathcal{K}_{\mathrm{z}}$ belong to a predefined function class such as affine, polynomial or rational, such that $\mathcal{B}_{\mathrm{z}}\circ\mu=\mathcal{B}$, $\mathcal{K}_{\mathrm{z}}\circ\mu=\mathcal{K}$ [Aut_Iacob_inputs], [Mohammadpour]. Note that the dependencies of $\mu$ are based on the choice of $\mathcal{B}$ and $\mathcal{K}$.
@@ 3.3 Identification problem
The objective is to introduce a parametrized version of (25) to learn the underlying dependencies together with a lifting map using ANNs. This means identifying the lifted state $z_{k}$, the linear maps $A$, $C$, and the nonlinear maps $B$ and $K$. To this end, we introduce an identification cost function that we chose to be the squared prediction error due to its extended use and success in system identification and its close connection to maximum-likelihood estimators under specific settings [book_ljung], [Ljung2013]. For this, a predictor is needed and we derive it next. As a first step, we exploit $e_{k}=y_{k}-Cz_{k}$ in the assumed innovation form (21) and substitute it in (25a) to obtain:
@@
with $\tilde{K}(z_{k},u_{k},y_{k})\mathrel{\mathop{\ordinarycolon}}=K(z_{k},u_{k},y_{k}-Cz_{k})$. Then, iterating (25b) forward in time, for $n\geq 1$, we arrive at
@@ 3.4 Subspace encoder
To overcome problem (i), in this section, the estimation of the lifted state $z_{k}$ is considered in terms of an encoder. By exploiting input-output data, we use the reconsutructability concept, discussed in the autonomous case, which we now generalize for (25). Starting with observability, the following equality holds based on (30):
@@
where, as in the autonomous case, we have the set of nonlinear constraints $\Psi$. For $n\geq 1$, if $\exists(z_{*},w_{*})\in\mathbb{R}^{n_{\mathrm{z}}}\times\mathbb{R}^{nn_{\mathrm{u}}\times(n+1)n_{\mathrm{y}}\times(n+1)n_{\mathrm{y}}}$ for which the Jacobian $\nabla_{(z_{\ast},w_{\ast})}\mathcal{O}_{\mathrm{z},n}$ has full row rank $n_{\mathrm{z}}$, then there exist open sets $\mathbb{Z}_{0}\subseteq\mathbb{R}^{n_{\mathrm{z}}}$, $\mathbb{U}_{0}\subseteq\mathbb{R}^{n_{\mathrm{u}}}$, $\mathbb{Y}_{0}\subseteq\mathbb{R}^{n_{\mathrm{y}}}$, $\mathbb{E}_{0}\subseteq\mathbb{R}^{n_{\mathrm{y}}}$, corresponding to the neighborhood of $(z_{*},w_{*})$ for which $\mathcal{O}_{\mathrm{z},n}$ is partially invertible and (25) is locally observable on $(\mathbb{Z}_{0},\mathbb{U}_{0},\mathbb{Y}_{0},\mathbb{E}_{0})$, see [nl_obs_1982]. Note that if the representation is locally observable, then the above condition is satisfied for any $n\geq n_{\mathrm{z}}-1$ By inverting $\mathcal{O}_{\mathrm{z},n}$, we get
@@
where $\Lambda_{\mathrm{z},n}\mathrel{\mathop{\ordinarycolon}}\mathbb{U}^{n}_{0}\times\mathbb{Y}^{n+1}_{0}\times\mathbb{E}^{n+1}_{0}\rightarrow{\mathbb{R}^{n_{\mathrm{z}}}}$ is the observability map. To determine $z_{k}$ based on past input-output data, we derive
@@
$\mathrel{\mathop{\ordinarycolon}}=\Pi_{\mathrm{z},n}(u^{k-1}_{k-n},y^{k}_{k-n},e^{k}_{k-n})$
where $\Pi_{\mathrm{z},n}\mathrel{\mathop{\ordinarycolon}}\mathbb{U}^{n}_{0}\times\mathbb{Y}^{n+1}_{0}\times\mathbb{E}^{n+1}_{0}\rightarrow{\mathbb{R}^{n_{\mathrm{z}}}}$ is the recosntructability map. Note that the noise sequence $e^{k}_{k-n}$ is not directly available in practice, hence to compute $z_{k}$ based on (36), again we can exploit the i.i.d. white noise property of $e_{k}$ to arrive at:
@@
which mapping gives an efficient estimator of $z_{k}$ in the conditional mean sense based on past data with a given lag $n$. In fact, (37) functions as an encoder, mapping from the past data to the lifted state $z_{k}$, i.e., a subspace of the lifted state space. However, an exact calculation of the encoder $\bar{\Pi}_{\mathrm{z},n}$ for a given ANN parametrization of $f_{\theta}$ and $h_{\theta}$ is infeasible in practice, due to the required analytic inversion of $\mathcal{O}_{\mathrm{z},n}$ to get $\Lambda_{\mathrm{z},n}$ and the computation of the conditional expectation of $\bar{\Pi}_{\mathrm{z},n}$ under the unknown probability density function of $e_{k}$. Hence, our objective is to learn $\bar{\Pi}_{\mathrm{z},n}$ directly from the data by introducing a parametrized function $\bar{\Pi}^{\eta}_{\mathrm{z},n}$ with parameters $\eta\in\Upsilon\subseteq\mathbb{R}^{n_{\eta}}$, e.g., using an ANN in the general case, which is co-estimated with $A_{\theta}$, $B_{\theta}$, $C_{\theta}$ and $K_{\theta}$ [^4] We note an alternate method described in [otto_bilinear], where the authors describe the Koopman representation as a hidden Markov model and use Kalman filtering to estimate the latent (lifted) state based on partial observations. However, the bilinear form and the way noise affects the lifted model are chosen empirically, whereas (32) represents an analytical form..
Note that, similar to the autonomous case discussed in Section 2.3, we can conceptually show that exploiting the observability and reconstructability properties of the nonlinear system (21), the potentially needed number of lags $n\geq n_{\mathrm{z}}-1$ is greatly reduced. Given a lifting map $\Phi\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{n_{\mathrm{x}}}\rightarrow\mathbb{R}^{n_{\mathrm{z}}}$, the state $z_{k}=\Phi(x_{k})$ can be calculated based on the reconstructed $x_{k}$ using a number of lags $n\geq n_{\mathrm{x}}-1$, as detailed in [Gerben_Aut], for the reconstructabillity map associated to (21). As such, when computing $\bar{\Pi}^{\eta}_{\mathrm{z},n}$, the number of lags needed to estimate $z_{k}$ is related to the dimension of the underlying nonlinear system, rather than that of the lifted system. It is important to note that, while $n_{\mathrm{x}}-1$ guarantees local invertibility, for global observability guarantees one would need to increase the number of lags $n$. For example, works such as [Stark_1] and [Stark_2] describe a sufficient condition for reconstruction of $x_{k}$ to be $n\geq 2n_{\mathrm{x}}$, for nonlinear systems with deterministic and stochastic forcing, respectively. Note that, as the last step in (29) is $k+n$ instead of k+n1k+n-1, we subtract 1 from the given value in [Stark_1] and [Stark_2], which is $2n_{\mathrm{x}}+1$. While (34) is more complex than the maps discussed in [Stark_1] and [Stark_2], these results can still serve as a guideline when performing experiments and choosing $n$. Furthermore, if $n_{\mathrm{x}}$ is known, although $\mathcal{O}_{\mathrm{z},n}$ may be locally invertible for smaller values of $n$, it is still a safe choice to start with $n\geq n_{\mathrm{x}}-1$, which provides local guarantees for reconstructability.